OpenStudy (anonymous):

a lab emp is mixing a 10% saline solution with 4% saline solution how much of each solution is needed to make 500 millilters of a 6% solution? a.170 mililiters of 10% solution and 340 milliters of 4% solution b.167 milliters of 10% solution and 333 milliter of 4% c.110 millilters of 10% solution and 210 milliters of 4% solutions d .155 milliters of the 10% solution and 300ml of 4% solution does anyone know ?

OpenStudy (anonymous):

10%*v1+4%*v2=total solute v1+v2=total volume=500 ml concentration = solute/volume=(.*v1+.04*v2)/(v1+v2)=.06

OpenStudy (anonymous):

i dont understand